# Nonasymptotic Convergence Rates for Plug-and-Play Methods With MMSE Denoisers

### v4

@@ Abstract
It is known that the minimum-mean-squared-error (MMSE) denoiser under Gaussian noise can be written as a proximal operator, which suffices for asymptotic convergence of plug-and-play (PnP) methods but does not reveal the structure of the induced regularizer or give convergence rates. We show that the MMSE denoiser corresponds to a regularizer that can be written explicitly as an upper Moreau envelope of the negative log-marginal density, which in turn implies that the regularizer is $1$-weakly convex. Using this property, we derive (to the best of our knowledge) the first sublinear convergence guarantee for PnP proximal gradient descent with an MMSE denoiser. We validate the theory with a one-dimensional synthetic study that recovers the implicit regularizer. We also validate the theory with imaging experiments (deblurring and computed tomography), which exhibit the predicted sublinear behavior.
I Background
In this work, we focus on the setting of linear inverse problems, where $x\in\mathbb{R}^{n}$ is the true signal, $y\in\mathbb{R}^{m}$ is the observed signal obtained by
@@ I Background
for linear forward operator $A\in\mathbb{R}^{m\times n}$ and measurement noise $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$. The task of recovering $x$ from $y$ is extensively studied in statistics and signal processing. A special case is the denoising problem, obtained when $A=I$:
@@
Recovering $x$ from $y$ is often posed as the minimization of a regularized objective function
@@
where $f$ is a data-fidelity term (typically $f(x)\coloneqq\tfrac{1}{2}\|Ax-y\|^{2}$ under Gaussian noise), and $g$ is a regularization term that encodes prior beliefs about the class of admissible solutions. For instance, the $\ell^{1}$-norm ($g(x)\propto\lVert x\rVert_{1}$) promotes sparsity and underlies the theory of compressed sensing [candes_robust, donoho_compressed, bora_compressed]. Similarly, total variation (TV), defined as $g(x)\propto\operatorname{TV}(x)\coloneqq\sum_{k}|x_{k}-x_{k-1}|$ favors piecewise-constant solutions and is widely used in image denoising [rudin_nonlinear, chambolle_algorithm, chambolle_first, liu_image, parhi2023sparsity]. More recently, much of research in imaging inverse problems has shifted towards learned regularizers [hertrich2025learning, ducotterd2025learning, ducotterd2025multivariate, goujon2023neural, pourya2025dealing, fang_whats, ongie2020deep, kamilov_plug]. Without regularization, the problem is often ill-posed in the sense of Hadamard [engl_regularization, hadamard_problemes]: Solutions may be non-unique, unstable, or nonexistent. Since the data-fidelity term is typically well-behaved, the regularizer largely determines the structure of solutions. More broadly, the regularizer largely dictates how optimization algorithms behave, so understanding its structure is crucial.
I-A Proximal Gradient Descent.
@@
$x_{k+1}\leftarrow x_{k}-\gamma\nabla F(x_{k}).$ (4)
With appropriately chosen step size $\gamma>0$, the iterates $x_{k}$ converge to a stationary point [kakade_dimension]. However, gradient descent is not well-suited for nonsmooth objectives.
@@
Here $\operatorname{prox}_{\gamma g}$ denotes the proximal operator of $g$ (see Section II), originally introduced by Moreau in [moreau_functions]. Convergence of PGD has been widely studied in both convex and non-convex settings and can be established under various assumptions on $f$ and $g$ with appropriately chosen step sizes $\gamma>0$ [li_accelerated, Table 1].
@@ I-B Plug-and-Play Methods
The plug-and-play (PnP) framework [venkatakrishnan_plug, kamilov_plug] extends PGD by replacing the proximal step $\operatorname{prox}_{\gamma g}$ with a generic denoiser $D_{\sigma}$ [boyd_distributed]. This substitution is motivated by the interpretation of $\operatorname{prox}_{\gamma g}$ as the maximum a posteriori (MAP) solution to the denoising problem (2) under prior $p_{X}\propto e^{-\gamma g}$ [combettes_proximal, Eq. (10.13)], the exact problem denoisers are designed to solve. The PnP iteration is given by
@@
and has been shown to yield state-of-the-art results when combined with powerful image denoisers such as BM3D [dabov_image] or a deep neural network (DNN) [jain_natural, zhang_beyond]. Often, we use data fidelity term $f(x)\coloneqq\tfrac{1}{2}\|Ax-y\|^{2}$, in which case the PnP iteration becomes
@@
A central challenge in the PnP framework is that while a proximal operator is always tied to a well-defined regularizer, many of the most effective denoising methods are not. This substitution therefore breaks the direct connection to an explicit optimization problem, making theoretical interpretations and convergence claims difficult. Much of the literature imposes strict assumptions on the denoiser to obtain convergence guarantees. Many of the most effective denoisers do not satisfy these assumptions, yet still achieve excellent results (for example [ahmad_plug]).
I-C Related Works.
@@
When $f$ and $g$ are convex with $L$-smooth $f$, it is well known that PGD converges to a fixed point for appropriate step size $\gamma$ [parikh_prox, Section 4.2]. This result can be proven via the machinery of monotone operator theory, the goal being to show that the composite PGD step (5) constitutes an averaged operator, guaranteeing convergence [combettes_signal, combettes_solving, bauschke_convex]. [^1] When $g$ is proper, lower-semicontinuous, convex, $\partial g$ is maximally monotone, making $\operatorname{prox}_{\gamma g}=(I+\gamma\partial g)^{-1}$ firmly-nonexpansive, i.e., $1/2$-averaged. The gradient step is also averaged for $L$-smooth $f$, making their composition an averaged operator.
@@
In particular, convergence of PnP-PGD has been shown under tight assumptions on the denoiser, such as averaged/nonexpansive behavior [sun_online] or suitable Lipschitz (e.g., residual-Lipschitz) bounds [ryu_plug]. Many works impose these assumptions through architectural constraints on the learned denoisers via, e.g., spectral normalization [ryu_plug] or through explicit training techniques [terris_building, bredies_learning]. In [pesquet_learning], the authors propose a technique for learning a maximally monotone operator. Convergence of PnP has also been shown for specific denoisers, such as kernel denoisers [gavaskar_plug], denoisers that can be written as the gradient step of a smooth function [hurault_proximal], bounded denoisers [chan_plug], and linear denoisers [nair_fixed].
@@
Instead of treating the denoiser as a black box with certain conditions, a complementary line of work leverages its intrinsic statistical structure. The MMSE estimator under Gaussian noise is a notable example. With the Tweedie/Stein identities [efron_tweedies, stein_bound], the MMSE denoiser can be tied to the score function, endowing it with useful structural properties [milanfar_denoising]. Further, [gribonval_reconciling] and [gribonval_should] showed that the MMSE denoiser is $C^{\infty}$ and can be written as the proximal map of a regularizer (which is $C^{\infty}$) under various forward and noise models, and derived an explicit formula for this regularizer (28). This progress enabled the derivation of asymptotic convergence for both the ADMM [park_convergence] and PGD [xu_provable] variants of PnP using an MMSE denoiser. However, a gap in the theory remains: The explicit formula offers little insight into the behavior of the induced regularizer, and there are no nonasymptotic convergence rates tailored to MMSE denoisers.
I-D Contributions.
@@
    $\sigma^{2}\sup_{z}\mathopen{}\mathclose{{\left(f_{Z}(z)-\tfrac{1}{2\sigma^{2}}\|z-x\|^{2}}}\right)$ (8)
    where $f_{Z}=-\log p_{Z}$ is the negative log marginal distribution in the noise model Eq. 2. This is an explicit and interpretable description of the regularizer implicit in the MMSE denoiser, and we believe that this identity is new.
  • Nonasymptotic convergence of MMSE-PGD via weak convexity. In Section III-A, we prove that $x\mapsto\phi_{\mathrm{MMSE}}(x)+\tfrac{1}{2}\|x\|^{2}$ is convex, hence $\phi_{\mathrm{MMSE}}$ is $1$-weakly convex. With this property we establish (to the best of our knowledge) the first nonasymptotic convergence guarantees for PGD-PnP with an MMSE denoiser in Theorem III.7. In particular, we show that, under an $L$-smooth data-fidelity term $f$ with $L<1$, PGD reaches an $\mathcal{O}(1/\sqrt{k})$-stationary point after $k$ iterations.
  • Experiments. In Section IV, we consider the following experimental setups: (i) A one-dimensional synthetic example validating the upper Moreau envelope form of the regularizer implicit in the MMSE denoiser by comparing a learned DNN estimator to its explicit form, (ii) Gaussian deblurring on the MNIST dataset [lecun_mnist], and (iii) computed tomography on the MayoCT dataset [mcollough_low]. In both imaging tasks we observe the predicted sublinear $\mathcal{O}(1/\sqrt{k}$) convergence in the subgradient residual.
@@ II Preliminaries
In this section, we collect some definitions and results used in the remainder of the paper. First, we recall the definition of the convex conjugate, also known as the Legendre-Fenchel transform or Fenchel conjugate. See for example [rockafellar_variational].
Definition II.1.
For a function $f:\mathbb{R}^{n}\to\mathbb{R}$, the convex conjugate $f^{*}$ is given by
@@
Definition II.2 (Lower Moreau Envelope).
$M_{\gamma}f(x)\coloneqq\inf_{y}\mathopen{}\mathclose{{\left(f(y)+\tfrac{1}{2\gamma}\|y-x\|^{2}}}\right).$ (10)
Definition II.3 (Upper Moreau Envelope).
$M^{\gamma}f(x)\coloneqq\sup_{y}\mathopen{}\mathclose{{\left(f(y)-\tfrac{1}{2\gamma}\|y-x\|^{2}}}\right).$ (11)
Note that the lower Moreau envelope is the usual Moreau envelope, but we make the distinction in this paper since we use both lower and upper Moreau envelopes. For the infimum defining the lower Moreau envelope to be achieved, $f$ must be lower-semicontinuous. An important property of the lower Moreau envelope $M_{\gamma}f$ is that it shares global minimizers with $f$ whenever $f$ is lower-semicontinuous, i.e. $\operatorname*{arg\,min}_{x}f(x)=\operatorname*{arg\,min}_{x}M_{\gamma}f(x).$ See [bernard_lasry] for further background.
Definition II.4 (Weak convexity).
Crucial to our analysis is the following lemma, where we include a short proof for completeness since the proof in the original reference had typos.
Lemma II.5 ([bernard_lasry, Lemma 3]).
If $f:\mathbb{R}^{n}\to\mathbb{R}$ is locally bounded, $M^{\gamma}M_{\gamma}f(x)\leq f(x)$, and equality holds if and only if $f$ is $\tfrac{1}{\gamma}$-weakly convex.
Proof.
From the definition,
$\phantom{{}={}}M^{\gamma}M_{\gamma}f(x)$
$=\sup_{y}\inf_{z}\mathopen{}\mathclose{{\left(f(z)+\tfrac{1}{2\gamma}\|z-y\|^{2}-\tfrac{1}{2\gamma}\|y-x\|^{2}}}\right)$
$\leq\sup_{y}f(x)=f(x),$ (12)
by taking $z=x$. We now write
$\phantom{{}={}}M^{\gamma}M_{\gamma}f(x)+\tfrac{1}{2\gamma}\|x\|^{2}$
@@
The convex conjugate $g^{*}$ evaluated at $y/\gamma$ can be written as
$g^{*}(y/\gamma)$ $=\sup_{z}\mathopen{}\mathclose{{\left(\langle y/\gamma,z\rangle-g(z)}}\right)$
$=-\inf_{z}\mathopen{}\mathclose{{\left(g(z)-\langle y/\gamma,z\rangle}}\right),$ (14)
so that (13) can be written as
$\phantom{{}={}}\sup_{y}\mathopen{}\mathclose{{\left(-g^{*}(y/\gamma)+\langle y/\gamma,x\rangle}}\right)$
$=\sup_{y}\mathopen{}\mathclose{{\left(\langle y,x\rangle-g^{*}(y)}}\right)=g^{**}(x).$ (15)
Therefore, we have
$M^{\gamma}M_{\gamma}f(x)+\tfrac{1}{2\gamma}\|x\|^{2}=\mathopen{}\mathclose{{\left(f+\tfrac{1}{2\gamma}\|\cdot\|^{2}}}\right)^{**}(x),$ (16)
and the result follows when we recall that a function equals its biconjugate if and only if it is convex [rockafellar_convex, Section 12.2]. ∎
We now turn to the proximal operator, a central tool in optimization. We will later show that MMSE estimators themselves admit a proximal representation.
Definition II.6 (Proximal operator).
Let $\gamma>0$, and $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be proper and lower-semicontinuous. The proximal operator of $f$ with parameter $\gamma$ is the mapping
[…]
Note that $\operatorname{prox}_{\gamma f}$ is often set-valued. By coercivity of the quadratic term, $\operatorname{prox}_{\gamma f}(x)$ is non-empty for every $x\in\mathbb{R}^{n}$. For a proper, lower-semicontinuous function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$, the infimum defining the lower Moreau envelope is attained at every minimizer of the proximal problem; that is,
[…]
for all $x\in\mathbb{R}^{n}$. When $f$ is a convex, lower-semicontinuous proper function, it is well known that $\operatorname{prox}_{\gamma f}$ and $M_{\gamma}f$ are also related by the following formula due to Moreau [moreau_proximite]:
@@
Although the proximal operator of a non-convex function is generally set-valued, when $\operatorname{prox}_{\gamma f}$ is an MMSE denoiser, it is always single-valued and continuous, which is the setting of our main results. This motivates the following extension of the Moreau gradient identity.
Theorem II.7 (Extension of the Moreau Gradient Identity).
Let $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be proper and lower-semicontinuous. If $\operatorname{prox}_{\gamma f}(x)$ is single-valued and continuous for all $x$, then the gradient of the lower Moreau envelope exists and takes the form
@@
From the definition, $M_{\gamma}f(x)=\min_{z}h(x,z)$. Fix some $\bar{x}\in\mathbb{R}^{n}$ where there is a unique minimizer $\{\bar{z}\}=\operatorname{prox}_{\gamma f}(\bar{x})$ by assumption. Take some $u\in\mathbb{R}^{n}$. There is some $\{z_{t}\}=\operatorname{prox}_{\gamma f}(\bar{x}+tu)$ so that
$\phantom{{}={}}M_{\gamma}f(\bar{x}+tu)=h(\bar{x}+tu,z_{t})$
$=h(\bar{x},z_{t})+\tfrac{t}{\gamma}\langle\bar{x}-z_{t},u\rangle+\tfrac{t^{2}}{2\gamma}\|u\|^{2}$
$\geq M_{\gamma}f(\bar{x})+\tfrac{t}{\gamma}\langle\bar{x}-z_{t},u\rangle+\tfrac{t^{2}}{2\gamma}\|u\|^{2}.$ (22)
@@
$\phantom{{}={}}\liminf_{t\to 0}\tfrac{M_{\gamma}f(\bar{x}+tu)-M_{\gamma}f(\bar{x})}{t}$
$\geq\liminf_{t\to 0}\mathopen{}\mathclose{{\left(\tfrac{1}{\gamma}\langle\bar{x}-z_{t},u\rangle+\tfrac{t}{2\gamma}\|u\|^{2}}}\right)$
$=\tfrac{1}{\gamma}\langle\bar{x}-\bar{z},u\rangle,$ (23)
@@
$\phantom{{}={}}M_{\gamma}f(\bar{x}+tu)\leq h(\bar{x}+tu,\bar{z})$
$=M_{\gamma}f(\bar{x})+\tfrac{t}{\gamma}\langle\bar{x}-\bar{z},u\rangle+\tfrac{t^{2}}{2\gamma}\|u\|^{2}.$ (24)
@@
$\phantom{{}={}}\limsup_{t\to 0}\tfrac{M_{\gamma}f(\bar{x}+tu)-M_{\gamma}f(\bar{x})}{t}$
$\leq\limsup_{t\to 0}\mathopen{}\mathclose{{\left(\tfrac{1}{\gamma}\langle\bar{x}-\bar{z},u\rangle+\tfrac{t}{2\gamma}\|u\|^{2}}}\right)$
$=\tfrac{1}{\gamma}\langle\bar{x}-\bar{z},u\rangle.$ (25)
Combining (II) and (II), we have that the directional derivative of $M_{\gamma}f$ at $\bar{x}$ in direction $u$ satisfies $\nabla_{u}M_{\gamma}f(\bar{x})=\tfrac{1}{\gamma}\langle\bar{x}-\bar{z},u\rangle$. Therefore, the gradient is
$\nabla M_{\gamma}f(x)=\tfrac{1}{\gamma}(\bar{x}-\bar{z})=\tfrac{1}{\gamma}\mathopen{}\mathclose{{\left(\bar{x}-\operatorname{prox}_{\gamma f}(\bar{x})}}\right).$ (26)
Many similar versions of this identity are well known in the literature. [^2] Results of this type are often formulated in terms of prox-regularity. We avoid this terminology here, as we don’t use it anywhere else and the proof of the stated result is quite simple. For more detail see [rockafellar_variational, Theorem 13.37].
@@
Theorem II.8 (Tweedie’s Formula).

### v6

@@ Abstract
It is known that the minimum-mean-squared-error (MMSE) denoiser under Gaussian noise can be written as a proximal operator, which suffices for asymptotic convergence of plug-and-play (PnP) methods but does not reveal the structure of the induced regularizer or give convergence rates. We show that the MMSE denoiser corresponds to a regularizer that can be written explicitly as an upper Moreau envelope of the negative log-marginal density, which in turn implies that the regularizer is 1-weakly convex on its domain. Using this property, we derive (to the best of our knowledge) the first sublinear convergence guarantee for PnP proximal gradient descent with an MMSE denoiser. We validate the theory with a one-dimensional synthetic study that recovers the implicit regularizer. We also validate the theory with imaging experiments (deblurring and computed tomography), which exhibit the predicted sublinear behavior.
I Introduction
In this work, we focus on the setting of linear inverse problems, where $\overline{x}\in\mathbb{R}^{n}$ is the true signal, $y\in\mathbb{R}^{m}$ is the observed signal obtained by
@@ I Introduction
for linear forward operator $A\in\mathbb{R}^{m\times n}$ and measurement noise $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$. The task of recovering $\overline{x}$ from $y$ is extensively studied in statistics and signal processing. A special case is the denoising problem, obtained when $A=I$:
@@
Recovering $\overline{x}$ from $y$ is often posed as the minimization of a regularized objective function
@@
where $f$ is a data-fidelity term (typically $f(x)\coloneqq\tfrac{1}{2}\|Ax-y\|^{2}$ under Gaussian noise), and $g$ is a regularization term that encodes prior beliefs about the class of admissible solutions. For instance, the $\ell^{1}$-norm ($g(x)\propto\lVert x\rVert_{1}$) promotes sparsity and underlies the theory of compressed sensing [11, 19, 8]. Similarly, total variation (TV), defined as $g(x)\propto\operatorname{TV}(x)\coloneqq\sum_{i}|x_{i}-x_{i-1}|$ for $1$-D signals, favors piecewise-constant solutions and is widely used in image denoising [57, 13, 12, 42, 51]. Without regularization, the problem is often ill-posed in the sense of Hadamard [23, 32]: Solutions may be non-unique, unstable, or nonexistent. Since the data-fidelity term is typically well-behaved, the regularizer largely determines the structure of solutions. More broadly, the regularizer largely dictates how optimization algorithms behave, so understanding its structure is crucial.
I-A Proximal Gradient Descent
@@
$x_{k+1}=x_{k}-\gamma\nabla F(x_{k}).$ (4)
With appropriately chosen step size $\gamma>0$, the iterates $x_{k}$ converge to a stationary point [38]. However, gradient descent is not well-suited for nonsmooth objectives.
@@
Here $\operatorname{prox}_{\gamma g}$ denotes the proximal operator of $g$ (see Section II), originally introduced by Moreau in [46]. Convergence of PGD has been widely studied in both convex and nonconvex settings and can be established under various assumptions on $f$ and $g$ with appropriately chosen step sizes $\gamma>0$ [41, Table 1].
@@ I-B Plug-and-Play Methods
The plug-and-play (PnP) framework [64, 39] extends PGD by replacing the proximal step $\operatorname{prox}_{\gamma g}$ with a generic denoiser $D_{\sigma}$ [9]. This substitution is motivated by the interpretation of $\operatorname{prox}_{\gamma g}$ as the maximum a posteriori (MAP) solution to the denoising problem (2) under prior $p_{X}\propto\mathrm{e}^{-\gamma g}$ [15, Eq. (10.13)], the exact problem denoisers are designed to solve. The PnP-PGD iteration is given by
@@
and has been shown to yield state-of-the-art results when combined with powerful image denoisers such as BM3D [18] or a deep neural network (DNN) [37, 66]. Often, we use data fidelity term $f(x)\coloneqq\tfrac{1}{2}\|Ax-y\|^{2}$, in which case the PnP iteration becomes
@@
A central challenge in the PnP framework is that while a proximal operator is always tied to a well-defined regularizer, many of the most effective denoising methods are not. This substitution therefore breaks the direct connection to an explicit optimization problem, making theoretical interpretations and convergence claims difficult. Much of the literature imposes strict assumptions on the denoiser to obtain convergence guarantees. Many of the most effective denoisers do not satisfy these assumptions, yet still achieve excellent results [2].
TABLE I: Summary of related convergence guarantees for plug-and-play proximal gradient descent (PnP-PGD). Note $L_{f}$ is the Lipschitz constant for $\nabla f$, $L_{g_{\sigma}}$ is the Lipschitz constant for $\nabla g_{\sigma}$, $\gamma$ is the step size, and $K$ is the iterate.
Work Assumptions on data-fidelity term $f$ Assumptions on denoiser $D_{\sigma}$ Convergence rate
Classical PGD [6] Convex, $L_{f}$-smooth $D_{\sigma}=\operatorname{prox}_{g}$, where $g$ is convex $\mathcal{O}(1/\sqrt{K})$
Lipschitz condition on the denoiser [58] $L_{f}$-smooth, $\mu$-Strongly convex $D_{\sigma}-I$ is $\varepsilon$-Lipschitz for $\varepsilon<\tfrac{2\mu}{L_{f}-\mu}$ Asymptotic
Gradient-step denoisers [34, 35, 36] $L_{f}$-smooth, $\gamma L_{f}<\tfrac{L_{g_{\sigma}}+2}{L_{g_{\sigma}}+1}$ $D_{\sigma}=I-\nabla g_{\sigma}$, where $L_{g_{\sigma}}<1$ $\mathcal{O}(1/\sqrt{K})$
Prior work with MMSE denoisers [65] $L_{f}$-smooth, $L_{f}<1$ MMSE denoiser Asymptotic
This paper (Section III-B) $L_{f}$-smooth, $L_{f}<1$ MMSE denoiser $\mathcal{O}(1/\sqrt{K})$
I-C Related Works
@@
When $f$ and $g$ are convex with $L$-smooth $f$, it is well known that PGD converges to a fixed point for appropriate step size $\gamma$ [52, Section 4.2]. This result can be proven via the machinery of monotone operator theory, the goal being to show that the composite PGD step (5) constitutes an averaged operator, guaranteeing convergence [16, 17, 5]. [^1] When $g$ is proper, lower-semicontinuous, convex, $\partial g$ is maximally monotone, making $\operatorname{prox}_{\gamma g}=(I+\gamma\partial g)^{-1}$ firmly-nonexpansive, i.e., $1/2$-averaged. The gradient step is also averaged for $L$-smooth $f$, making their composition an averaged operator.
@@
In particular, convergence of PnP-PGD has been shown under tight assumptions on the denoiser, such as nonexpansiveness [61] or suitable Lipschitz (e.g., residual-Lipschitz) bounds [58]. Many works impose these assumptions through architectural constraints on the learned denoisers via, e.g., spectral normalization [58] or through explicit training techniques [62, 10]. In [54], the authors propose a technique for learning a maximally monotone operator. Convergence of PnP has also been shown for specific denoisers, such as kernel denoisers [26, 3], gradient-step (GS) denoisers [36, 34, 35], bounded denoisers [14], and linear denoisers [48, 25]. A recent line of work [36] has significantly relaxed the assumptions on GS denoisers. We note that this setting (see [36, Proposition 3.1]) still does not apply to the MMSE denoiser, as it cannot be written as the gradient of a smooth (Lipschitz gradient) function.
A similar line of work learns regularizers explicitly [33, 20, 21, 27, 55, 24, 49, 39, 59, 28, 50], rather than implicitly through the denoiser. This approach benefits from direct interpretability of the regularizer and can inherit convergence guarantees from classical variational theory.
@@
A complementary approach leverages the intrinsic statistical structure of the denoiser. A particularly notable example—and the subject of this work—is the MMSE denoiser under Gaussian noise. With the Tweedie/Stein identities [22, 60], the MMSE denoiser can be tied to the score function, endowing it with useful structural properties [45]. Further, [29] and [31] showed that the MMSE denoiser is $C^{\infty}$ and can be written as the proximal map of a regularizer (which is $C^{\infty}$) under various forward and noise models, and derived an explicit formula for this regularizer (24). This progress enabled the derivation of asymptotic convergence for both the ADMM [53] and PGD [65] variants of PnP using an MMSE denoiser. However, a gap in the theory remains: The explicit formula offers little insight into the behavior of the induced regularizer, and there are no nonasymptotic convergence rates tailored to MMSE denoisers. Table I highlights several closely related results on PnP-PGD, the assumptions they impose on the denoiser, and the convergence guarantees they establish.
I-D Contributions
@@
    $\sigma^{2}\sup_{z}\mathopen{}\mathclose{{\left\{f_{Z}(z)-\tfrac{1}{2\sigma^{2}}\|z-x\|^{2}}}\right\},$ (8)
    on its domain, where $f_{Z}=-\log p_{Z}$ is the negative log marginal distribution from Eq. 2. This is a novel explicit form of the implicit regularizer in the MMSE denoiser.
  • Nonasymptotic convergence of MMSE-PGD via weak convexity. In Theorem III.2, we prove that $\phi_{\mathrm{MMSE}}$ is $1$-weakly convex on its domain. With this property we establish (to the best of our knowledge) the first nonasymptotic convergence guarantees for PGD-PnP with an MMSE denoiser in Theorem III.3 and Section III-B. In particular, we show that, under an $L$-smooth data-fidelity term $f$ with $L<1$, both $\|\nabla F(x_{k})\|$ and $\|x_{k+1}-x_{k}\|$ decrease sublinearly in terms of the best iterate.
  • Experiments. In Section IV, we consider the following experimental setups: (i) A one-dimensional synthetic example validating the upper Moreau envelope form of the implicit regularizer in the MMSE denoiser by comparing a learned DNN denoiser to its explicit form, (ii) Gaussian deblurring on the MNIST dataset [40], and (iii) computed tomography on the MayoCT dataset [44]. In both imaging tasks we observe the predicted sublinear convergence.
@@ II Preliminaries
In this section, we collect some definitions and results used in the remainder of the paper.
Definition II.1 (Lower Moreau Envelope).
$M_{\gamma}f(x)\coloneqq\inf_{y}\mathopen{}\mathclose{{\left\{f(y)+\tfrac{1}{2\gamma}\|y-x\|^{2}}}\right\}.$ (9)
Definition II.2 (Upper Moreau Envelope).
$M^{\gamma}f(x)\coloneqq\sup_{y}\mathopen{}\mathclose{{\left\{f(y)-\tfrac{1}{2\gamma}\|y-x\|^{2}}}\right\}.$ (10)
Note that the lower Moreau envelope is the usual Moreau envelope, but we make the distinction in this paper since we use both lower and upper Moreau envelopes. For the infimum defining the lower Moreau envelope to be achieved, $f$ must be lower-semicontinuous. An important property of the lower Moreau envelope $M_{\gamma}f$ is that it shares global minimizers with $f$ whenever $f$ is lower-semicontinuous, i.e. $\operatorname*{arg\,min}_{x}f(x)=\operatorname*{arg\,min}_{x}M_{\gamma}f(x).$ See [7] for further background.
Definition II.3 (Weak convexity).
Crucial to our analysis is the following lemma, adapted from [7, Lemma 3].
Lemma II.4.
Let $f:\mathbb{R}^{n}\to(-\infty,\infty]$. Then, $M^{\gamma}M_{\gamma}f(x)\leq f(x)$ for all $x\in\mathbb{R}^{n}$ and $M^{\gamma}f$ is $\tfrac{1}{\gamma}$-weakly convex.
Proof.
By definition,
$M^{\gamma}M_{\gamma}f(x)$
$={}$ $\sup_{y}\inf_{z}\mathopen{}\mathclose{{\left\{f(z)+\tfrac{1}{2\gamma}\|z-y\|^{2}-\tfrac{1}{2\gamma}\|y-x\|^{2}}}\right\}$
$\leq{}$ $\sup_{y}f(x)=f(x).$ (11)
Next,
$M^{\gamma}f(x)+\tfrac{1}{2\gamma}\|x\|^{2}$
$={}$ $\sup_{y}\mathopen{}\mathclose{{\left\{f(y)-\tfrac{1}{2\gamma}\|x-y\|^{2}+\tfrac{1}{2\gamma}\|x\|^{2}}}\right\}$
$={}$ $\sup_{y}\mathopen{}\mathclose{{\left\{f(y)-\tfrac{1}{2\gamma}\|y\|^{2}+\tfrac{1}{\gamma}\langle x,y\rangle}}\right\},$ (12)
which is a supremum of affine functions, and hence convex. Therefore, $M^{\gamma}f$ is $\tfrac{1}{\gamma}$-weakly convex. ∎
We now turn to the proximal operator, a central tool in optimization. We shall rely on the fact that MMSE denoisers themselves admit a proximal representation.
Definition II.5 (Proximal operator).
Let $\gamma>0$, and $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$ be proper and lower-semicontinuous. The proximal operator of $f$ with parameter $\gamma$ is the mapping
@@
Note that $\operatorname{prox}_{\gamma f}$ is often set-valued. By coercivity of the quadratic term, $\operatorname{prox}_{\gamma f}(x)$ is non-empty for every $x\in\mathbb{R}^{n}$. For a proper, lower-semicontinuous function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$, the infimum defining the lower Moreau envelope is attained at every minimizer of the proximal problem; that is,
@@
for all $x\in\mathbb{R}^{n}$. When $f$ is a convex, lower-semicontinuous, and proper function, it is well known that $\operatorname{prox}_{\gamma f}$ and $M_{\gamma}f$ are also related by the following formula due to Moreau [47]:
@@
Although the proximal operator of a nonconvex function is generally set-valued, when $\operatorname{prox}_{\gamma f}$ is an MMSE denoiser, it is always single-valued and continuous, which is the setting of our main results. This motivates the following extension of the Moreau gradient identity.
Theorem II.6 (Extension of the Moreau Gradient Identity).
Let $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$ be proper and lower-semicontinuous. If $\operatorname{prox}_{\gamma f}(x)$ is single-valued and continuous for all $x$, then the gradient of the lower Moreau envelope exists and takes the form
@@
From the definition, $M_{\gamma}f(x)=\min_{z}h(x,z)$. Fix some $x_{0}\in\mathbb{R}^{n}$ where there is a unique minimizer $\{z_{0}\}=\operatorname{prox}_{\gamma f}(x_{0})$ by assumption. Take some $u\in\mathbb{R}^{n}$. There is some $\{z_{t}\}=\operatorname{prox}_{\gamma f}(x_{0}+tu)$ so that
$M_{\gamma}f(x_{0}+tu)$
$={}$ $h(x_{0}+tu,z_{t})$
$={}$ $h(x_{0},z_{t})+\tfrac{t}{\gamma}\langle x_{0}-z_{t},u\rangle+\tfrac{t^{2}}{2\gamma}\|u\|^{2}$
$\geq{}$ $M_{\gamma}f(x_{0})+\tfrac{t}{\gamma}\langle x_{0}-z_{t},u\rangle+\tfrac{t^{2}}{2\gamma}\|u\|^{2}.$ (18)
@@
$\liminf_{t\to 0}\tfrac{M_{\gamma}f(x_{0}+tu)-M_{\gamma}f(x_{0})}{t}$
$\geq{}$ $\liminf_{t\to 0}\mathopen{}\mathclose{{\left(\tfrac{1}{\gamma}\langle x_{0}-z_{t},u\rangle+\tfrac{t}{2\gamma}\|u\|^{2}}}\right)$
$={}$ $\tfrac{1}{\gamma}\langle x_{0}-z_{0},u\rangle,$ (19)
@@
$M_{\gamma}f(x_{0}+tu)\leq h(x_{0}+tu,z_{0})$
$={}$ $M_{\gamma}f(x_{0})+\tfrac{t}{\gamma}\langle x_{0}-z_{0},u\rangle+\tfrac{t^{2}}{2\gamma}\|u\|^{2}.$ (20)
@@
$\limsup_{t\to 0}\tfrac{M_{\gamma}f(x_{0}+tu)-M_{\gamma}f(x_{0})}{t}$
$\leq{}$ $\limsup_{t\to 0}\mathopen{}\mathclose{{\left(\tfrac{1}{\gamma}\langle x_{0}-z_{0},u\rangle+\tfrac{t}{2\gamma}\|u\|^{2}}}\right)$
$={}$ $\tfrac{1}{\gamma}\langle x_{0}-z_{0},u\rangle.$ (21)
Combining (II) and (II), we have that the directional derivative of $M_{\gamma}f$ at $x_{0}$ in direction $u$ satisfies $(M_{\gamma}f)^{\prime}(x_{0},u)=\tfrac{1}{\gamma}\langle x_{0}-z_{0},u\rangle$. Therefore, the gradient is
$\nabla M_{\gamma}f(x_{0})=\tfrac{1}{\gamma}(x_{0}-z_{0})=\tfrac{1}{\gamma}\mathopen{}\mathclose{{\left(x_{0}-\operatorname{prox}_{\gamma f}(x_{0})}}\right),$ (22)
which completes the proof. ∎
Many similar versions of this identity are well known in the literature, but are often formulated in terms of prox-regularity. We included a simple self-contained proof for completeness. For more detail, see [56, Theorem 13.37].
@@
Theorem II.7 (Tweedie’s Formula).